If 3 digits are chosen and each digit is equally likely to be chosen, what is the probability that all digits are the same?
Answer by venugopalramana(3286)

If 3 digits are chosen and each digit is equally likely to be chosen, what is the probability that all digits are the same?
THERE ARE 10 DIGITS INCLUDING 0
TOTAL NUMBER OF ALTERNATIVES POSSIBLE
I DIGIT = 10
II DIGIT =10
III DIGIT=10
TOTAL NUMBER = 10*10*10 =1000
TAKING SAME DIGIT
I DIGIT = 10 WAYS
II DIGIT =1 WAY
III DIGIT =1 WAY
TOTAL NUMBER = 10*1*1=10
PROBABILITY = 10/1000=1/100=0.01